OpenStudy (anonymous):

*help requested* *fan and medal awarded* Determine the type of boundary line and shading for the graph of the inequality -3x + y > 6. Dashed line with shading on the side that includes the origin. Solid line with shading on the side that does not include the origin. Dashed line with shading on the side that does not include the origin. Solid line with shading on the side that includes the origin A cell phone plan offers a monthly rate of $25 for unlimited minutes plus $.05 per text message. Which algebraic expression below represents the total cost for this cell phone plan in term

OpenStudy (mathstudent55):

To graph an inequality of this type, first look at the corresponding equation, 3x + y = 6 Graph it using intercepts. To do that, let x = 0, then 3(0) + y = 6; y = 6. That's the first point, (0, 6) now let y = 0, 3x + 0 = 6; 3x = 6; x = 2. There's your second point, (2, 0) Since the inequality has a less than sign, not less than or equal to, use a dashed line to graph the line. Now pick 1 point on either side of the dashed line. If that point makes the original inequality true, shade the side of the line the point is in. If not, shade the other side.

OpenStudy (mathstudent55):